OpenStudy (anonymous):

what is x=6x-40 some one please help i need this to pass my mid term exam

OpenStudy (texaschic101):

x = 6x - 40 --- subtract 6x from both sides x - 6x = -40 -5x = -40 -- divide both sides by -5 x = -40/-5 x = 8

OpenStudy (anonymous):

x=6x-40 -6x on both sides -5x=-40 /-5 on both sides -40/-5=8

OpenStudy (texaschic101):

great minds think alike :)

OpenStudy (anonymous):

its is for angle measure i got that same answer but it cant be 8 degrees

OpenStudy (texaschic101):

that is how to solve the equation. Is there more to the problem...since this is obviously a geometry problem

OpenStudy (anonymous):

yea it says find m<ABC

OpenStudy (texaschic101):

you got a picture ?

OpenStudy (texaschic101):

sorry...I kinda lost track of time...I have to go, but I will be back on later...again, I am sorry

OpenStudy (anonymous):

i can take one on my phone but i cantsend it on here

OpenStudy (anonymous):

nvm all i had to do was plug in the 8 to the original equation thank you for your help